OpenStudy (anonymous):

Check answer please? "Julie runs 3 miles less than twice the number of miles, n, that Hannah runs". 2n - 3 3n - 2<

OpenStudy (shamil98):

3 miles less, 3 is subtracted. twice the number, n so 2n - 3

OpenStudy (anonymous):

Ok thanks can you check two more please?

OpenStudy (anonymous):

Solve 5s + 40 = 55 using number sense 3 15 19 95

OpenStudy (anonymous):

I think 19 would be the answer.

OpenStudy (shamil98):

What do you think it is? 5s + 40 = 55 - 40 -40 5s = 15

OpenStudy (shamil98):

divide both sides by 5 to get s.

OpenStudy (shamil98):

why do you think its 19?

OpenStudy (anonymous):

Ok got it thanks again last one.

OpenStudy (anonymous):

Translate " the sum of 8 and the product of 2 and y is 24 " in to an equation. 8y+2=24 8y-2=24 8+2y=24 8-2y=24

OpenStudy (anonymous):

I say C.

OpenStudy (shamil98):

correct

OpenStudy (anonymous):

Yes, Thanks